OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles. C = 30°, a = 28, c = 14

OpenStudy (anonymous):

options: 1 0 2

OpenStudy (anonymous):

@.Sam. @A.Avinash_Goutham @amistre64 @blues @experimentX @ParthKohli @phi

OpenStudy (phi):

can you use the Law of Sines ?

OpenStudy (anonymous):

im a bit confused by it

OpenStudy (phi):

write down sin angle over length of the side opposite that angle. they tell you angle C and side c (normally people label the sides opposite the angle with a lower case letter.... in other words side c is opposite angle C) so the first step is write the ratio: \[ \frac{\sin C}{c} \] but put in the numbers

OpenStudy (phi):

now make another ratio. they give you side a, but not angle A you can write the ratio \[ \frac{\sin A}{28} \] set the two ratios equal to each other can you do that ?

OpenStudy (anonymous):

can you show me, it would help more

OpenStudy (phi):

first, can you replace the letters with their numbers (from the question) \[ \frac{\sin C}{c} \]